# How Do Recommendation Models Amplify Popularity Bias? An Analysis from the Spectral Perspective

### v3

@@ 5.1 Experiment Settings
For fair comparisons, we implement all compared methods with uniform MF backbone and MSE loss. We also explore the performance with other backbones and losses in subsection 5.3. Besides above baselines, we also compare our method with the methods on mitigating dimension collapse, including nCL [9] and DirectAU [53]; and the debiasing methods tailored for GNN-based methods including APDA [73] and GCFlogdet [69] when using GNN-based backbones.
Parameter Settings. The embedding dimension $d$ is 256 while other dimensions are explored in B.4 . Grid search is utilized find the optimal hyperparameters. More details refer to Appendix C.2.
5.2 Performance Comparison (RQ1)
Comparison under three testing paradigms. Table 2 showcases the NDCG@K comparison across seven datasets over three testing paradigms. Under the Common testing paradigm, our ReSN, with few exceptions, consistently outperforms compared methods. This superior performance can be attributed to the rigorous theoretical foundations of ReSN, which pinpoint and address the root cause of bias amplification. By curbing this bias amplification, ReSN achieves significant improvements in recommendation accuracy. Transitioning to the Debiased and Uniform-exposure testing paradigms, the improvements by ReSN become even more impressive, demonstrating its effectiveness in mitigating popularity bias.
@@ 5.2 Performance Comparison (RQ1)
5.3 Adaptability Exploration (RQ2)
Table 3: NDCG@20 comparison with LightGCN and XSimGCL backbone.
@@ 5.3 Adaptability Exploration (RQ2)
To investigate the adaptability of ReSN, we evaluate it with various backbones and loss functions under debiased testing paradigm. Table 3 showcases the performance of ReSN under LightGCN [27] and XSimGCL [62], where APDA [73] and GCFlogdet [69] that tailored for GNN-based backbones are included. Besides, Table 6 depicts the results with BCE and BPR losses. Notably, ReSN consistently outperforms compared methods, irrespective of the chosen backbone or loss function. These findings affirm the great adaptability of ReSN, underscoring its ability to seamlessly integrate with diverse recommendation models. Our ReSN also outperforms those debiasing methods tailored for GNN-based methods. They reason is that they only consider to mitigate bias amplification raised by GNNs while ignoring the bias from the generic recommendation mechanism.
5.4 Hyperparameter Study (RQ3)
Figures 6 presents the recommendation accuracy, the ratio of popular items, and the ratio of principle singular value ($\sigma_{1}^{2}/\sum_{1\leq k\leq L}{\sigma_{k}^{2}}$) in ReSN as the hyperparameter $\beta$ varies. Notably, as $\beta$ increases, the ratio of principle singular value and the severity of popularity bias reduces. This trend affirms the efficacy of our regularizer. Regarding recommendation accuracy, it initially rises and then declines with an increase in $\beta$. This can be attributed to the fact that popularity bias isn’t intrinsically detrimental [68, 70]. Indeed, item popularity can also convey beneficial information about the item’s appeal or quality, which ought to be retained. Hence, the strategic approach for popularity bias lies in mitigating its bias amplification, rather than eliminating it entirely. That is also our target.
Figure 5: NDCG@20 comparison with different Loss functions.
@@ 5.4 Hyperparameter Study (RQ3)
Figure 6: The proportion of popular items in recommendations and the ratio of the largest singular value ($\sigma_{1}^{2}/\sum_{1\leq k\leq L}{\sigma_{k}^{2}}$) and NDCG@20 with varying $\beta$.
6 Related Work
@@
Analyses on Popularity Bias. In RS, items frequently exhibit a long-tailed distribution in terms of interaction frequency. Models trained on skewed data are susceptible to inheriting and exacerbating such bias [28, 74, 4, 5, 75]. The crux of tackling popularity bias lies in understanding why and how recommendation models intensify popularity bias. Several recent efforts aim to elucidate this. Among these, causality-based investigations stand out. For instance, Zhang et al. [68] developed a causal graph of the data generative process, attributing the amplification of popularity bias to a confounding effect; Wei et al. [56] presented an alternate causal graph, exploring the direct and indirect causal influence of popularity bias on predictions; Wang et al. [55] used yet another causal graph to highlight bias amplification, focusing on the impact of users’ historical long-tailed distribution across item groups. Despite their contributions, a common limitation among these causality-based methods is their surface-level engagement with the causal relationships among variables, rather than delving deeper into the underlying mechanisms. For example, these studies usually operate on the assumption that item popularity directly affects predictions. However, the specifics of how and why predictions memorize and are influenced by item popularity remain largely unexplored. Worse still, their effectiveness hinges on the accuracy of their respective causal graphs, which might not always align with real-world scenarios due to the widespread presence of unmeasured confounders [21, 33].
@@
There have been other theoretical investigations into popularity bias. For instance, Zhu et al. [74] demonstrate that model predictions inherit item popularity, yet they failed to elucidate the amplification. Also, their conclusions rely on a strong assumption that the preference scores maintain same distribution over different user-item pairs. The study by [41] shed light on the limited expressiveness of low-rank embeddings, giving clues of popularity bias in recommendations. Yet, they did not factor in the impact of long-tailed training data. In fact, popularity bias origins from long-tailed data [68, 74], amplified during training, which would be more serious than the theoretically analyses presented in [41]. Some efforts Chen et al. [13], Kim et al. [30] examined popularity bias through embedding magnitude, their theoretical analysis can only applied in the early stages of training. Other researchers delved into how graph neural networks amplify popularity bias, whether through influence functions [10], the hub effect [73] or dimensional collapse [69]. However, their conclusions can not be extended to general recommendation models.
@@
Methods on Tackling Popularity Bias. Recent efforts on addressing popularity bias are mainly four types: 1) Causality-driven methods assume a causal graph to identify popularity bias and employ causal inference techniques for rectification. While they have demonstrated efficacy, their success is closely tied to the accuracy of the causal graph. This poses challenges due to the prevalence of unmeasured confounders [21, 33, 40]. 2) Propensity-based methods [49, 25, 19, 66, 54] adjust the data distribution by reweighting the training data instances. While this approach directly negates popularity bias in the data, it may inadvertently obscure other valuable signals, such as item quality. Consequently, these methods often underperform compared to causality-driven ones. 3) Regularizer-based methods [3, 74, 47, 36, 29] constrain predictions by introducing regularization terms. For example, Zhu et al. [74] employs a Pearson coefficient regularizer to diminish the correlation between item popularity and model predictions; Rhee et al. [47] proposes to regularize the score differences; [36] constrains the predictions with IPL criterion. As discussed in section 4.2, their constraints are too strong, may significantly compromising accuracy. 4) Disentanglement-based methods [64, 59, 14] target at learning disentangled embeddings that segregate the influence of popularity from genuine user preferences. While promising, achieving a perfect disentanglement of popularity bias from true preferences remains a formidable challenge in RS.
7 Conclusion and Limitation
@@
In this study, we delve into the root cause of popularity bias amplification. Our analyses offer two core insights: 1) Item popularity is encoded in the principal spectrum of model predictions; 2) The phenomenon of dimension collapse accentuates the influence of the principal spectrum. Based on these insights, we introduce ReSN, an efficient technique aimed at mitigating popularity bias by penalizing the principle singular value. A potential limitation of our study pertains to the static perspective on popularity bias, neglecting its dynamic nature as it evolves temporally. It could be more insightful to investigate the mechanism of bias amplification in the context of temporal sequential recommendations, and to examine its evolution during the feedback loop.
@@ Appendix A Theoretical Analysis
A.1 Analysis of BPR loss
A.1.1 Reconceptualizing from the BPR to the BCE loss
BPR loss is expressed as:
${\mathcal{L}_{BPR}}=-\sum\limits_{u\in\mathcal{U}}{\sum\limits_{i\in\mathcal{I},{y_{ui}}=1}{\sum\limits_{j\in\mathcal{I},{y_{uj}}=0}\log({\mu({\bf{u}}_{u}^{\top}{{\bf{v}}_{i}}-{\bf{u}}_{u}^{\top}{{\bf{v}}_{j}})})}}$ (6)
Then we can construct a hyper-item space denoted as $\mathcal{I}^{\prime}=\mathcal{I}\times\mathcal{I}$ derived from item pairs, and define the embeddings of hyper-items as ${\mathbf{v}^{\prime}_{ij}}={{\mathbf{v}}_{i}}-{{\mathbf{v}}_{j}}$ and assign new observed interactions to the combinations of users and hyper-items, i.e., ${y^{\prime}_{u,ij}}=1$ for $y_{ui}=1\&y_{uj}=0$ and ${y^{\prime}_{u,ij}}=0$ for $y_{ui}=0\&y_{uj}=1$.
With this transformation, the BPR loss can be effectively reframed as:
${\mathcal{L}_{BPR}}$ $=-\sum\limits_{u\in\mathcal{U}}{\sum\limits_{i\in\mathcal{I},{y_{ui}}=1}{\sum\limits_{j\in\mathcal{I},{y_{uj}}=0}\log({\mu({\bf{u}}_{u}^{\top}{{\bf{v}}_{i}}-{\bf{u}}_{u}^{\top}{{\bf{v}}_{j}})})}}$
$=-\sum\limits_{u\in\mathcal{U}}\sum\limits_{i\in\mathcal{I},{y_{ui}}=1}\sum\limits_{j\in\mathcal{I},{y_{uj}}=0}\log(\mu({\bf{u}}_{u}^{\top}({\bf{v}}_{i}-{\bf{v}}_{j})))$
$=-\frac{1}{2}\sum\limits_{u\in\mathcal{U}}\left(\sum\limits_{i\in\mathcal{I},{y_{ui}}=1}{\sum\limits_{j\in\mathcal{I},{y_{uj}}=0}}\log(\mu({\bf{u}}_{u}^{\top}({\bf{v}}_{i}-{\bf{v}}_{j})))+\sum\limits_{i\in\mathcal{I},{y_{ui}}=0}{\sum\limits_{j\in\mathcal{I},{y_{uj}}=1}}\log(\mu({\bf{u}}_{u}^{\top}({\bf{v}}_{j}-{\bf{v}}_{i})))\right)$
$=-\frac{1}{2}\sum\limits_{u\in\mathcal{U}}\left(\sum_{(i,j)\in\mathcal{I^{\prime}}y^{\prime}_{u,ij}=1}\log(\mu({\bf{u}}_{u}^{\top}{\mathbf{v}^{\prime}_{ij}}))+\sum_{(i,j)\in\mathcal{I^{\prime}}y^{\prime}_{u,ij}=0}\log(\mu(-{\bf{u}}_{u}^{\top}{\mathbf{v}^{\prime}_{ij}}))\right)$
If the activate function $\mu(.)$ is taken as the Sigmoid function $\sigma(.)$, we have:
${\mathcal{L}_{BPR}}$ $=-\frac{1}{2}\sum\limits_{u\in\mathcal{U}}\left(\sum_{(i,j)\in\mathcal{I^{\prime}}y^{\prime}_{u,ij}=1}\log(\sigma({\bf{u}}_{u}^{\top}{\mathbf{v}^{\prime}_{ij}}))+\sum_{(i,j)\in\mathcal{I^{\prime}}y^{\prime}_{u,ij}=0}\log(1-\sigma({\bf{u}}_{u}^{\top}{\mathbf{v}^{\prime}_{ij}}))\right)$ (7)
$=-\frac{1}{2}\sum\limits_{u\in\mathcal{U}}\sum_{(i,j)\in\mathcal{I^{\prime}}}\left(y^{\prime}_{u,ij}\log(\sigma({\bf{u}}_{u}^{\top}{\mathbf{v}^{\prime}_{ij}}))+(1-y^{\prime}_{u,ij})\log(1-\sigma({\bf{u}}_{u}^{\top}{\mathbf{v}^{\prime}_{ij}}))\right)$
From Equation 7, the BPR loss can be reframed as the BCE loss with the hyper-items space $\mathcal{I^{\prime}}$. This allows the analyses on point-wise losses to be easily extended to BPR loss.
A.1.2 Suitability of ReSN for BPR Loss
As delineated above, BPR can be regarded as a specialized point-wise loss by introducing the concept of hyper-items. Let’s denote the embeddings of these hyper-items with matrix $\mathbf{V}^{\prime}\in\mathbb{R}^{m^{2}\times d}$, where the ($i\cdot m+j$)-th row corresponds to the embedding of the $ij$-th hyper-item ${\mathbf{v}^{\prime}_{ij}}={{\mathbf{v}}_{i}}-{{\mathbf{v}}_{j}}$. One might infer that an apt regularizer for the BPR loss would be $\frac{||{\bf{U}}{{\bf{V^{\prime}}}^{\top}}{\bf{V^{\prime}}}{{\bf{U}}^{\top}}{\bf{e}}||^{2}}{||{\bf{V^{\prime}}}{{\bf{U}}^{\top}}{\bf{e}}||^{2}}$, rather than $\frac{||{\bf{U}}{{\bf{V}}^{\top}}{\bf{V}}{{\bf{U}}^{\top}}{\bf{e}}||^{2}}{||{\bf{V}}{{\bf{U}}^{\top}}{\bf{e}}||^{2}}$.
As for the numerator, we have:
${{{\bf{V^{\prime}}}}^{\top}}{\bf{V^{\prime}}}$ $=\sum\limits_{i,j\in I}{({{\bf{v}}_{i}}-{{\bf{v}}_{j}})^{\top}{{({{\bf{v}}_{i}}-{{\bf{v}}_{j}})}}}$ (8)
$=2m\sum\limits_{i\in I}{{{\bf{v}}_{i}}^{\top}{\bf{v}}_{i}}-2(\sum\limits_{i\in I}{{{\bf{v}}_{i}}})^{\top}{(\sum\limits_{i\in I}{{{\bf{v}}_{i}}})}=2m\mathbf{V}^{\top}{\mathbf{V}}-2m^{2}{\bf{\bar{v}}}^{\top}{{{\bf{\bar{v}}}}}$
where $\bar{\mathbf{v}}=\sum_{i\in I}{{{\bf{v}}_{i}}}/m$ denote the mean vector of the item embeddings. Furthermore, current literature posits that an ideal item representation should emulate a uniform distribution over the unit ball [53]. This implies that $\bar{\mathbf{v}}$ tends to gravitate towards the origin. Thus, ${{{\bf{V^{\prime}}}}^{\top}}{\bf{V^{\prime}}}$ can be approximated by $\mathbf{V}^{\top}{\mathbf{V}}$ and $||{\bf{U}}{{\bf{V}^{\prime}}^{\top}}{\bf{V}^{\prime}}{{\bf{U}}^{\top}}{\bf{e}}||^{2}$ can be approximated by $||{\bf{U}}{{\bf{V}}^{\top}}{\bf{V}}{{\bf{U}}^{\top}}{\bf{e}}||^{2}$.
Similarly, for the denominator:
$||{\bf{V^{\prime}U^{\top}e}}||^{2}$ $=\sum\limits_{i,j\in I}{({{\bf{v}}_{i}}{\bf{U^{\top}e}}-{{\bf{v}}_{j}}{\bf{U^{\top}e}})^{\top}{{({{\bf{v}}_{i}}{\bf{U^{\top}e}}-{{\bf{v}}_{j}}{\bf{U^{\top}e}})}}}$ (9)
$=2m\sum\limits_{i\in I}({{\bf{v}}_{i}}{\bf{U^{\top}e}})^{\top}({{\bf{v}}_{i}}{\bf{U^{\top}e}})-2(\sum\limits_{i\in I}{{\bf{v}}_{i}}{\bf{U^{\top}e}})^{\top}(\sum\limits_{i\in I}{{\bf{v}}_{i}}{\bf{U^{\top}e}})$
$=2m({\bf{VU^{\top}e}})^{\top}({\bf{VU^{\top}e}})-2m^{2}({{\bf{\bar{v}}}}{\bf{U^{\top}e}})^{\top}({{\bf{\bar{v}}}}{\bf{U^{\top}e}})$
We can deduce ${||{\bf{V}^{\prime}}{{\bf{U}}^{\top}}{\bf{e}}||^{2}}$ can be approximted by ${||{\bf{V}}{{\bf{U}}^{\top}}{\bf{e}}||^{2}}$. Consequently, ReSN emerges as a logical regularizer even for the BPR loss. This assertion is also validated by our experiments.
A.2 Proof of Theorem 1

### v6

@@ 5.1. Experiment Settings
For fair comparisons, we implement all compared methods with uniform MF backbone and MSE loss. We also explore the performance with other backbones and losses in subsection 5.3. Besides above baselines, we also compare our method with the methods on mitigating dimension collapse, including nCL (Chen et al., 2023b) and DirectAU (Wang et al., 2022); and the debiasing methods tailored for GNN-based methods including APDA (Zhou et al., 2023) and GCFlogdet (Zhang et al., 2023b) when using GNN-based backbones.
Parameter Settings. The embedding dimension $d$ is 256 while other dimensions are explored in B.3 . Grid search is utilized find the optimal hyperparameters. More details refer to Appendix C.2
5.2. Performance Comparison (RQ1)
Comparison under three testing paradigms. Table 3 showcases the NDCG@20 comparison across seven datasets over three testing paradigms. Under the Common testing paradigm, our ReSN, with few exceptions, consistently outperforms compared methods. This superior performance can be attributed to the rigorous theoretical foundations of ReSN, which pinpoint and address the root cause of bias amplification. By curbing this bias amplification, ReSN achieves significant improvements in recommendation accuracy. Transitioning to the Debiased and Uniform-exposure testing paradigms, the improvements by ReSN become even more impressive, demonstrating its effectiveness in mitigating popularity bias.
Figure 3. Pareto curves of compared methods illustrating the trade-off between accuracy and fairness under the common testing paradigm.
@@ 5.2. Performance Comparison (RQ1)
5.3. Adaptability Exploration (RQ2)
Table 5. NDCG@20 comparison with GNN-based backbones (LightGCN, XSimGCL) under the debiased testing paradigm.
@@ 5.3. Adaptability Exploration (RQ2)
Table 6. NDCG@20 comparison with different Loss functions under the debiased testing paradigm.
@@
Figure 4. The proportion of popular items in recommendations and the ratio of the largest singular value ($\sigma_{1}^{2}/\sum_{1\leq k\leq L}{\sigma_{k}^{2}}$) and NDCG@20 with varying $\beta$.
Table 7. Running time comparison (s/Epoch), where ReSN-Direct directly calculates spectral norm without acceleration.
Movielens Douban Gowalla
MF 0.177 2.098 0.634
ReSN 0.181 2.124 0.675
ReSN-Direct 649 59239 15280
Speedup Ratio 3585 27890 22637
To investigate the adaptability of ReSN, we evaluate it with various backbones and loss functions under debiased testing paradigm. Table 5 showcases the performance of ReSN under LightGCN (He et al., 2020) and XSimGCL (Yu et al., 2023), where APDA (Zhou et al., 2023) and GCFlogdet (Zhang et al., 2023b) that tailored for GNN-based backbones are included. Besides, Table 6 depicts the results with BCE and BPR losses. Notably, ReSN consistently outperforms compared methods, irrespective of the chosen backbone or loss function. These findings affirm the great adaptability of ReSN, underscoring its ability to seamlessly integrate with diverse recommendation models. Our ReSN also outperforms those debiasing methods tailored for GNN-based methods. They reason is that they only consider to mitigate bias amplification raised by GNNs while ignoring the bias from the generic recommendation mechanism.
5.4. Hyperparameter Study (RQ3)
Figures 4 presents the recommendation accuracy, the ratio of popular items, and the ratio of principle singular value ($\sigma_{1}^{2}/\sum_{1\leq k\leq L}{\sigma_{k}^{2}}$) in ReSN as the hyperparameter $\beta$ varies. Notably, as $\beta$ increases, the ratio of principle singular value and the severity of popularity bias reduces. This trend affirms the efficacy of our regularizer. Regarding recommendation accuracy, it initially rises and then declines with an increase in $\beta$. This can be attributed to the fact that popularity bias isn’t intrinsically detrimental (Zhang et al., 2021; Zhao et al., 2022). Indeed, item popularity can also convey beneficial information about the item’s appeal or quality, which ought to be retained. Hence, the strategic approach for popularity bias lies in mitigating its bias amplification, rather than eliminating it entirely. That is also our target.
5.5. Efficiency Study (RQ4)
To further validate the effectiveness of our acceleration strategy, we test the running time per epoch of ReSN and the original brute-force strategy employed to compute the gradient of the spectral norm. Also, we present the baseline MF for comparison. The results are presented in Table 7. As can be seen, our acceleration strategy achieves over 3600, 27000 and 22000 times impressive speed-up in both datasets, respectively. Moreover, compared with MF, our ReSN does not incur much computational overhead.
6. Related Work
@@
Analyses on Popularity Bias. In RS, items frequently exhibit a long-tailed distribution in terms of interaction frequency. Models trained on skewed data are susceptible to inheriting and exacerbating such bias (Jannach et al., 2015; Zhu et al., 2021; Abdollahpouri et al., 2019, 2020; Zhu et al., 2022; Wu et al., 2024). The crux of tackling popularity bias lies in understanding why and how recommendation models intensify popularity bias. Several recent efforts aim to elucidate this. Among these, causality-based investigations stand out. For instance, Zhang et al. (2021); Wang et al. (2021) developed a causal graph of the data generative process, attributing the amplification of popularity bias to a confounding effect; Wei et al. (2021) presented an alternate causal graph, exploring the direct and indirect causal influence of popularity bias on predictions. A common limitation among these causality-based methods is their surface-level engagement with the causal relationships among variables, rather than delving deeper into the underlying mechanisms. For example, these studies usually operate on the assumption that item popularity directly affects predictions. However, the specifics of how and why predictions memorize and are influenced by item popularity remain largely unexplored. Worse still, their effectiveness hinges on the accuracy of their respective causal graphs, which might not always hold due to the unmeasured confounders (Ding et al., 2022; Li et al., 2023).
@@
There were other investigations into popularity bias. For instance, Zhu et al. (2021) demonstrate that model predictions inherit item popularity, yet they failed to elucidate the amplification. Also, their conclusions rely on a strong assumption that the preference scores maintain same distribution across different user-item pairs. The study by (Ohsaka and Togashi, 2023) shed light on the limited expressiveness of low-rank embeddings, giving clues of popularity bias in recommendations. Yet they did not factor in the impact of long-tailed training data. In fact, popularity bias origins from long-tailed data (Zhang et al., 2021; Zhu et al., 2021), amplified during training, which would be more serious than the theoretically analyses presented in (Ohsaka and Togashi, 2023). Some efforts (Chen et al., 2023d; Kim et al., 2023) examined popularity bias through embedding magnitude, their theoretical analysis can only applied in the early stages of training. Other researchers delved into how graph neural networks amplify popularity bias through influence functions (Chen et al., 2023c), the hub effect (Zhou et al., 2023) or dimensional collapse (Zhang et al., 2023b). However, their conclusions can not be extended to general recommendation models.
@@
Methods on Tackling Popularity Bias. Recent efforts on addressing popularity bias are mainly four types: 1) Causality-driven methods assume a causal graph to identify popularity bias and employ causal inference techniques for rectification. While they have demonstrated efficacy, their success is closely tied to the accuracy of the causal graph. This poses challenges due to the prevalence of unmeasured confounders (Ding et al., 2022; Li et al., 2023; Ning et al., 2024). 2) Propensity-based methods (Schnabel et al., 2016; Gruson et al., 2019; Chen et al., 2021; Zhang and Shen, 2023; Wang et al., 2024b) adjust the data distribution by reweighting the training data instances. While this approach directly negates popularity bias in the data, it may inadvertently obscure other valuable signals, such as item quality. Consequently, these methods often underperform compared to causality-driven ones. 3) Regularizer-based methods (Abdollahpouri et al., 2017; Zhu et al., 2021; Rhee et al., 2022; Liu et al., 2023; Jin et al., 2024) constrain predictions by introducing regularization terms. For example, Zhu et al. (2021) employs a Pearson coefficient regularizer to diminish the correlation between item popularity and model predictions; Zhang et al. (2023b) adopts a regularizer for mitigating embeddings collapse; Rhee et al. (2022) proposes to regularize the score differences; (Liu et al., 2023) constrains the predictions with IPL criterion. As discussed in section 2, their constraints are too strong, may significantly compromising accuracy. 4) Disentanglement-based methods (Zhang et al., 2023a; Xv et al., 2022; Chen et al., 2022) target at learning disentangled embeddings that segregate the influence of popularity from genuine user preferences. While promising, achieving a perfect disentanglement of popularity bias from true preferences remains a formidable challenge in RS.
Among the related work, the one most closely related to ours is (Zhang et al., 2023b), but we emphasize that our work differs in two key aspects: 1) Their theoretical justification of bias amplification focuses solely on GNNs, whereas our analysis applies to generic recommendation mechanisms. 2) Their regularizer aims to mitigate collapse of user/item embeddings, while our ReSN specifically targets the mitigation of the principal spectrum’s influence. Section 4.2 provides a detailed discussion of these differences, demonstrating that ReSN is more effective in debiasing. Table 4 also offers empirical evidence supporting our claims.
7. Conclusion
@@
In this study, we delve into the root cause of popularity bias amplification. Our analyses offer two core insights: 1) Item popularity is encoded in the principal spectrum of model predictions; 2) The phenomenon of dimension reduction accentuates the influence of the principal spectrum. Based on these insights, we introduce ReSN, an efficient technique aimed at mitigating popularity bias by penalizing the principle singular value. A potential limitation of our study pertains to the static perspective on popularity bias, neglecting its dynamic nature as it evolves temporally. It could be more insightful to investigate the mechanism of bias amplification in the context of temporal sequential recommendations, and to examine its evolution during the feedback loop.
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (62372399, 62476244) and the advanced computing resources provided by the Supercomputing Center of Hangzhou City University.
@@ Appendix A Theoretical Analysis
A.1. Proof of Theorem 1